OpenStudy (anonymous):

Medal for help please ty :) What is a equation in slope intercept form for the line that passes through the points (1,-3) and (3,1)? A.y=3x+1 B.y=x-3 c.y=2x+5 d.y=2x-5

OpenStudy (anonymous):

use this.. \[\frac{ y-y1 }{ x-x1 }=\frac{ y2-y1 }{ x2-x1 }\] plug values and get result..

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

Anyone know if i'm correct ?

OpenStudy (anonymous):

@MarrBear

OpenStudy (anonymous):

its D... y=2x-5

OpenStudy (anonymous):

D ^-^

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

yw